Integration by Partial Fractions


1.) ∫dt/(1+sint+cost)


2.) ∫dy/(5+4cosy)


1
Expert's answer
2021-05-07T10:20:30-0400

"1) \displaystyle\int\frac{dt}{1+\sin{t}+\cos{t}}
\\
2) \displaystyle\int\frac{dy}{5+4\cos{y}}"

Solution:

"1) \displaystyle\int\frac{dt}{1+\sin{t}+\cos{t}}"

Let's apply the change of variables:

"\tan{\frac t2}=x"; "\displaystyle\frac{dt}{2\cos^2{\frac t2}}=dx" ; "\displaystyle\frac12(1+\tan^2{\frac t2})dt=dx" ; "\displaystyle dt=\frac{2dx}{1+x^2}" .

"\displaystyle\sin{t}=\frac{2\tan{\frac t2}}{1+\tan^2{\frac t2}}=\frac{2x}{1+x^2}" ; "\displaystyle\cos{t}=\frac{1-\tan^2{\frac t2}}{1+\tan^2{\frac t2}}=\frac{1-x^2}{1+x^2}" .

"\displaystyle\int\frac{dt}{1+\sin{t}+\cos{t}}=" "\displaystyle\int\frac{2dx}{(1+x^2)(1+\frac{2x}{1+x^2}+\frac{1-x^2}{1+x^2})}=" "\displaystyle\int\frac{2dx}{(2+2x)}=" "\displaystyle\int\frac{dx}{(1+x)}=\ln|1+x|+C=\ln{\left|1+\tan{\frac t2}\right|}+C" .

(C is constant).

"2) \displaystyle\int\frac{dy}{5+4\cos{y}}"

Let's apply the change of variables:

"\tan{\frac y2}=x"; "\displaystyle\frac{dy}{2\cos^2{\frac y2}}=dx" ; "\displaystyle\frac12(1+\tan^2{\frac y2})dy=dx" ; "\displaystyle dy=\frac{2dx}{1+x^2}" .

"\displaystyle\cos{y}=\frac{1-\tan^2{\frac y2}}{1+\tan^2{\frac y2}}=\frac{1-x^2}{1+x^2}" .

"\displaystyle\int\frac{dy}{5+4\cos{y}}=\int\frac{2dx}{(1+x^2)(5+4\cdot\frac{1-x^2}{1+x^2})}=" "\displaystyle\int\frac{2dx}{9+x^2}=\frac23\int\frac{\frac{dx}{3}}{1+\frac{x^2}{9}}=" "\displaystyle\frac23\arctan{\frac{x}{3}}+C=\frac23\arctan{\left(\frac{1}{3}\tan{\frac{y}{2}}\right)}+C" .

(C is constant).

Answer:

1)"\displaystyle\int\frac{dt}{1+\sin{t}+\cos{t}}=\ln{\left|1+\tan{\frac t2}\right|}+C" ;

"2) \displaystyle\int\frac{dy}{5+4\cos{y}}=\frac23\arctan{\left(\frac{1}{3}\tan{\frac{y}{2}}\right)}+C" .